OpenStudy (anonymous):

A bookstore manager plans to spend $60 on pens and pencils. The cost of each pen is $0.50 and the cost of each pencil is $0.10. The equation 0.10x + 0.50y = 60 can be used to model this situation. Determine the slope-intercept form of this equation

OpenStudy (matt101):

All you need to do is isolate y (so you have the equation in y = mx + b form) and you're done.

OpenStudy (anonymous):

so 0.10x=60 ?

OpenStudy (matt101):

What happened to y? You need to rearrange the equation so that you have y = something.

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

y=0.50+0.10x=60 ?

OpenStudy (matt101):

Lol but now another = has appeared... To isolate y, first move 0.10x to the other side, so you get 0.50y = 60 - 0.10x. But you're looking for y by itself, not 0.50y. What would you have to do to change this?

OpenStudy (anonymous):

would u just take it out?

OpenStudy (matt101):

Take out 0.50 completely? That would change the equation! y = 60 - 0.10x is a different equation from 0.50y = 60 - 0.10x You would have to divide both sides by 0.5 (or multiply both by 2) to isolate y. This would give you y = 120 - 0.20x

OpenStudy (anonymous):

ohh okay i get it now thanks

OpenStudy (anonymous):

@matt101 nvm i dnt get it, i put that answer but it said it was wrong

OpenStudy (matt101):

Well they might want it as y = -0.20x + 120